OpenStudy (anonymous):

If the student hears a sound at 25 dB, what is the intensity of the sound?

OpenStudy (anonymous):

\[D=10\log_{10}\frac{ I }{ I _{T}} \] \[I _{T}=1.0 *10^{^{-12}}\] I= intensity of sound in W/m^2 D=decibel level of the sound